a stone propelled from a catapult with a speed of 50m/s attains a height of 100m calculate the angle of projection
Answer #1
What is its initial vertical speed Vo ?
v = Vo - 9.8 t
v = 0 at top
so
t = Vo/9.8

then height
h = Vo t - (9.8/2) t^2
100 = Vo^2/9.8 - (9.8/2) (Vo/9.8)^2

100 = Vo^2/(2*9.8)
Vo = 44.3 = speed sin theta
sin theta = 44.3/50
theta = 62.4 deg up from horizontal
answered by: Chelcie
Answer #2
Indicate your subject in the "School Subject" box, so those with expertise in the area will respond to the question
answered by: mom